OpenStudy (anonymous):

I have a question for a polynomial equation project. Suppose you know the volume of your object and the relation of the length to the width and height, but did not know the length. Rewrite the formula using the variable x for the length. Substitute the value of the volume for V and express the width and height of the object in terms of x plus or minus a constant. For example, if the height measures 4 inches longer than the length the expression would be (x+4). Then simplify the expression and write in standard form. My measurements are length: 11 Height: 4 width: 7 and volume: 308.

OpenStudy (amistre64):

let x = 11, be unknown to start with, height = 4 = 11+n = x + n width = 7 = 11+k = x+k volume = x(x+n)(x+k) = 308

OpenStudy (anonymous):

So that would be the equation?

OpenStudy (amistre64):

yes, but i left it up to you to solve for n and k, and to expand it into the form that it is asking for

OpenStudy (anonymous):

okay that helped thank you.

OpenStudy (amistre64):

good luck ;)